# GTA: A Benchmark for General Tool Agents

### v1

@@ Abstract
Significant focus has been placed on integrating large language models (LLMs) with various tools in developing general-purpose agents. This poses a challenge to LLMs’ tool-use capabilities. However, there are evident gaps between existing tool-use evaluations and real-world scenarios. Current evaluations often use AI-generated queries, single-step tasks, dummy tools, and text-only interactions, failing to reveal the agents’ real-world problem-solving abilities effectively. To address this, we propose GTA, a benchmark for General Tool Agents, featuring three main aspects: (i) Real user queries: human-written queries with simple real-world objectives but implicit tool-use, requiring the LLM to reason the suitable tools and plan the solution steps. (ii) Real deployed tools: an evaluation platform equipped with tools across perception, operation, logic, and creativity categories to evaluate the agents’ actual task execution performance. (iii) Real multimodal inputs: authentic image files, such as spatial scenes, web page screenshots, tables, code snippets, and printed/handwritten materials, used as the query contexts to align with real-world scenarios closely. We design 229 real-world tasks and executable tool chains to evaluate mainstream LLMs. Our findings show that real-world user queries are challenging for existing LLMs, with GPT-4 completing less than 50% of the tasks and most LLMs achieving below 25%. This evaluation reveals the bottlenecks in the tool-use capabilities of current LLMs in real-world scenarios, which provides future direction for advancing general-purpose tool agents.
@@ 1 Introduction
Despite the progress on benchmarking the tool-use capability of LLMs made by recent works, especially on collecting massive APIs and AI-generated user queries to enable scalable testing, there remain noticeable gaps regarding real-world scenarios, as shown in Table 1. First, AI-generated user queries, limited by the generative model, often result in overly brief or monotonous solutions. This is unsuitable for evaluating the reasoning and planning capability of agent systems, as shown in Table 2. Second, existing tool-use benchmarks mainly focus on text-formed user-agent interaction, lacking assessment of multimodal capabilities, thus falling short of aligning with real-world scenarios effectively. Third, existing tool-use evaluation approaches build up virtual tools. They can only evaluate isolated steps in the tool invocation chains, thus unable to reflect the agents’ capability to end-to-end accomplish complex tasks.
Figure 1: Some samples in the GTA benchmark. The user queries are human-designed, step-implicit, and settled in real-world scenarios. Multimodal context inputs are provided. Solving these queries is helpful for users, and complex for a LLM-based tool agent. The agent needs to use a combination of executable tools in perception, operation, logic, and creativity categories.
@@
    Real user queries. The user queries are designed by humans, rather than generated by AI, to reflect real-world tasks accurately. These queries describe tasks with clear objectives, but the tool-use steps are implicit. Thus, the LLM must employ reasoning to deduce the suitable tools required to address the given tasks. In this way, we avoid the drawbacks of using AI-generated queries in which the tool invocation steps are often explicitly hinted at. Moreover, each query requires multiple steps to resolve, necessitating the model to plan the sequence of tool invocations.
@@
    Real deployed tools. We provide an evaluation platform deployed with tools across various categories, such as perception, operation, logic, and creativity. All tools are executable rather than simulated by text description. For each task, a detailed and executable ground truth tool chain is provided, including each tool-use step and the final answer. Each step includes the tool name, argument value, and the tool return value. The detailed tool chains enable a fine-grained evaluation of the actual problem-solving abilities of tool agents.
@@
Table 2: Comparison of GTA queries with AI-generated queries. The steps and tool types for queries in ToolBench and m&m’s are explicitly stated, as marked in red and blue. The queries in APIBench are simple, only containing one step. Our GTA’s queries are both step-implicit and tool-implicit.
@@
We manually design 229 real-world tasks and corresponding executable tool chains to evaluate mainstream LLMs. We build a platform covering a total of 14 tools across perception, operation, logic, and creation categories. Tools and some data samples are illustrated in Figure 1. We design fine-grained tool evaluation metrics that cover the entire process of tool invocation. Our findings indicate that real-world scenario queries present challenges to existing LLMs, with GPT-4 completing fewer than 50% of the tasks and the majority of LLMs managing less than 25%.
@@
  2. An evaluation platform equipped with a rich variety of executable tools covering the categories of perception, operation, logic, and creativity. Fine-grained metrics are designed for tool-use, unveiling the reasoning and planning capabilities of tool-augmented LLMs in real-world scenarios.
@@ 2 Related Work
Tool-use evaluations. With the rise of LLM-based agents, many studies have been conducted to evaluate the tool-use capabilities of LLMs. ToolBench [22] collects RESTful APIs and leverages ChatGPT [1] to design tool-use tasks and corresponding tool chains. Two metrics, Pass Rate and Win Rate, are devised to evaluate the efficacy of tool use. APIBench [20] is a comprehensive dataset that includes APIs from HuggingFace, TorchHub, and TensorHub, with evaluation metrics focusing on Abstract Syntax Tree (AST) accuracy. API-Bank [11] comprises 53 commonly utilized APIs, such as SearchEngine, PlayMusic, BookHotel, and ImageCaption, along with a comprehensive tool-augmented LLM workflow, to evaluate the API calling, retrieving, and planning abilities. m&m’s [13] is a benchmark to evaluate tool-use for multi-step multimodal tasks. It aims to evaluate different planning strategies for LLMs as planning agents. Most of the aforementioned benchmarks, however, rely on AI-generated queries. The tool-use steps are explicitly and rigidly included. Thus these queries do not accurately represent real-world scenarios. Among many previous studies, GAIA [15] is renowned for its real-world scenario based benchmark aiming at evaluating general AI assistants. It evaluates more general results grounded in real-world interactions, which is closer to our work. It designs questions that are conceptually simple for humans yet challenging for most advanced AIs. However, GAIA focuses on artificial general intelligence (AGI). In contrast, GTA is designed to evaluate tool agents specifically, offering real-deployed tools and executable tool chains for a fine-grained evaluation in real-world scenarios.
@@ 3.1 Dataset Formulation
Given a set of tools $\mathcal{T}_{c}=\left\{t_{k}\right\}_{k=1}^{N}$, a sample in GTA is composed of five parts $(\mathcal{F},\mathcal{Q},\mathcal{T},\mathcal{C},\mathcal{A})$. Among these parts, $\mathcal{F}$ is a set of files containing one or two images. $\mathcal{Q}$ is a query based on $\mathcal{F}$. It is a real-world scenario based problem of simple form but needs to be solved through multiple steps with tools in $\mathcal{T}_{c}$. Which tools need to be used, and in what steps, are not explicitly included in the query. They require reasoning and planning by the LLM, which serves as a central controller. This procedure is given in the reference tool chain $\mathcal{C}=\{s_{i}\}_{i=1}^{m}$. The tool chain contains $m$ steps. Each step is $s_{i}=(t_{i},a_{i},r_{i})$, where $t_{i}$ is the tool used in step $i$. $a_{i}$ and $r_{i}$ indicate arguments and return values. $\mathcal{T}=\bigcup_{j=1}^{m}\{t_{j}\}\subseteq\mathcal{T}_{c}$ notes the set of tools involved in this query. $\mathcal{A}$ is the final answer yielded by the LLM after reasoning with tools.
In our setting, $\mathcal{T}_{c}$ contains 14 tools across four categories, including perception, operation, logic, and creativity. The full list of tools is shown in Figure 1, and more detailed information can be found in Appendix B.1. The queries $\mathcal{Q}$ are classified into three types: subjective, objective, and image generation. Examples of the three types of queries are shown in Appendix B.2. For a subjective query $\mathcal{Q}_{s}$, the final answer $\mathcal{A}$ is usually some descriptive text. It is not unique, but the general idea is the same. In this case, $\mathcal{A}$ contains a list of three reference answers. For an objective query $\mathcal{Q}_{o}$, $\mathcal{A}$ is a uniquely determined number or phrase. For an image generation query $\mathcal{Q}_{g}$, we only measure the tool call procedure and do not evaluate the generated image. In this situation, $\mathcal{A}=\emptyset$.
@@ 3.2 Query Construction
To construct $(\mathcal{F},\mathcal{Q},\mathcal{T})$, we first gather human-designed queries that meet three main principles: i) Given $\mathcal{T}\subseteq\mathcal{T}_{c}$, the task $(\mathcal{F},\mathcal{Q})$ can be solved with the capabilities enabled by tools in $\mathcal{T}$. ii) To evaluate LLMs’ reasoning and planning abilities, the tool invocation steps should not be explicitly stated in the queries. iii) The queries are meaningful and based on real-world scenarios. Satisfying all the principles simultaneously is challenging. It requires $\mathcal{F},\mathcal{Q}$, and $\mathcal{T}$ to match each other in a sensible and logical way. We use a query construction pipeline based on exemplar expansion, as shown in the first part of Figure 2. We first give some initial exemplars with diverse scenarios and tool combinations. Then we instruct annotators to create more queries based on the exemplars.
@@
Exemplar designed by experts. We first design some initial questions as exemplars, which are provided in Appendix C.1. These example questions are of diverse scenarios and contain different tool combinations. Every sample should comprise six components: $\mathcal{F}$ (image files), $\mathcal{Q}$ (queries), $\mathcal{T}$ (involved tools), $\mathcal{S}$ (solution steps), $\mathcal{A}$ (answers), and $\mathcal{E}$ (evidence). Image files $\mathcal{F}$ could be obtained from the internet and their URLs must be recorded. $\mathcal{F}$ could also be a photo taken or a diagram drawn by the annotators. The query $\mathcal{Q}$ needs to avoid obvious references to a specific tool. For example, the query please describe the image for me is unqualified since it obviously refers to the tool ImageDescription. The components $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{E}$, will not appear in the final dataset but are utilized to assist annotators in meeting the annotation requirements. $\mathcal{S}$ represents the steps required to solve the problem. Annotators should note down the steps, ensuring their number exceeds two. The answer $\mathcal{A}$ of objective queries should be given to guarantee there is a unique answer. To ensure the uniqueness, the answer should not be dependent on the images generated in previous steps. For example, the question what kind of animal is in the picture should not be asked after generate an image of an animal, as the answer is uncertain. For queries utilizing the GoogleSearch tool, $\mathcal{E}$ should include the answer’s URL and a screenshot pinpointing the answer’s location to verify the query’s searchability with the tool.
@@
Diversified expansion by annotators. After the initial exemplars are given, we instruct annotators to create more samples based on each exemplar. We adopt a diversified expansion strategy for the annotators to expand the questions based on the exemplars. The general idea is to keep the tool set $\mathcal{T}$ of the template unchanged or slightly modify it. Then annotators brainstorm scenarios different from the template. Further information on the diversified expansion approach is detailed in Appendix C.2. For each sample, we have crafted a manual expansion example to serve as guidance for the annotators. After the expansion process, we perform a quality check and manually filter out the questions that do not satisfy the expansion requirements. The instruction documents for annotators are reported in Appendix C.3.
@@
Considerations for the search tool. Search results may change over time making the web search tool special in the evaluation dataset. For queries requiring the GoogleSearch tool, we perform two constraints. Firstly, the question must necessitate the use of GoogleSearch, rather than relying on an LLM’s internal knowledge. This can be achieved by designing time-sensitive questions, such as what is the 2023 QS ranking of Tsinghua University rather than general inquiries like where is Tsinghua University located in China. We may also direct the query to a specific information source, for instance asking, what is the recipe for Ma Po Tofu according to the BBC Good Food website instead of a broad question like what is the recipe for Ma Po Tofu. Secondly, within an evaluation dataset, it is crucial to ensure that answers remain constant over time. We can specify a time frame, web page, or organization within the question to fulfill this criterion. An example would be what is the 2024 QS ranking of Tsinghua University, rather than what is the QS ranking of Tsinghua University.
@@ 3.3 Tool Chain Construction
Based on the $(\mathcal{F},\mathcal{Q},\mathcal{T})$ samples constructed in Section 3.2, we instruct three annotators majoring in computer science to manually construct the corresponding tool chain $\mathcal{C}$ and the final answer $\mathcal{A}$. We design a JSON file structure, containing the query-related tool list, image paths, and ReAct [30] style dialog sequences. The dialog sequences include the user query, the executable tool chain, and the final answer. Initially, $(\mathcal{T},\mathcal{F},\mathcal{Q})$ are put into the associated sections for tools, images, and user queries. Subsequently, we deploy all tools in $\mathcal{T}_{c}$. The annotators utilize the tools according to the reference steps $\mathcal{S}$ and get the outcomes. They record this process in the tool chain section of the dialog sequences, alongside the final answer. Since we do not evaluate the tools’ efficacy, when a tool fails to provide accurate recognition for a query (for instance, OCR inaccuracies in text recognition within diagrams), we discard the query. Through the above process, we ensure the feasibility of the questions, the executability of the tool chains, as well as the precision of the final answers. The structure of the tool chain is provided in Appendix C.4.
@@
Total tool calls $728$

### v2

@@ Abstract
Significant focus has been placed on integrating large language models (LLMs) with various tools in developing general-purpose agents. This poses a challenge to LLMs’ tool-use capabilities. However, there are evident gaps between existing tool-use evaluations and real-world scenarios. Current evaluations often use AI-generated queries, single-step tasks, dummy tools, and text-only interactions, failing to effectively reveal the agents’ real-world problem-solving abilities. To address this, we propose GTA, a benchmark for General Tool Agents, featuring three main aspects: (i) Real user queries: human-written queries with simple real-world objectives but implicit tool-use, requiring the LLM to reason the suitable tools and plan the solution steps. (ii) Real deployed tools: an evaluation platform equipped with tools across perception, operation, logic, and creativity categories to evaluate the agents’ actual task execution performance. (iii) Real multimodal inputs: authentic image files, such as spatial scenes, web page screenshots, tables, code snippets, and printed/handwritten materials, used as the query contexts to align with real-world scenarios closely. We design 229 real-world tasks and executable tool chains to evaluate mainstream LLMs. Our findings show that real-world user queries are challenging for existing LLMs, with GPT-4 completing less than 50% of the tasks and most LLMs achieving below 25%. This evaluation reveals the bottlenecks in the tool-use capabilities of current LLMs in real-world scenarios, which provides future direction for advancing general-purpose tool agents. Dataset and code are available at https://github.com/open-compass/GTA.
@@ 1 Introduction
Despite the progress on benchmarking the tool-use capability of LLMs made by recent works, especially on collecting massive APIs and AI-generated user queries to enable scalable testing, there remain noticeable gaps regarding real-world scenarios, as shown in Table 1. First, AI-generated user queries, limited by the generative model, often result in overly brief or monotonous solutions. This is unsuitable for evaluating agent systems’ reasoning and planning capability, as shown in Table 2. Second, existing tool-use benchmarks mainly focus on text-formed user-agent interaction, lacking assessment of multimodal capabilities, thus falling short of aligning with real-world scenarios effectively. Third, existing tool-use evaluation approaches build up virtual tools. They can only evaluate isolated steps in the tool invocation chains, thus unable to reflect the agents’ capability to accomplish complex tasks end-to-end.
Figure 1: Some samples in the GTA benchmark. The user queries are human-designed, step-implicit, tool-implicit, and settled in real-world scenarios. Multimodal context inputs are provided. Solving these queries is helpful for users and complex for a LLM-based tool agent. The agent must use a combination of executable tools in perception, operation, logic, and creativity categories.
@@
    Real user queries. The user queries are designed by humans, rather than generated by AI, to reflect real-world tasks accurately. These queries describe tasks with clear objectives, but the tool-use steps are implicit. Thus, the LLM must use reasoning to deduce the suitable tools to address the given tasks. In this way, we avoid the drawbacks of using AI-generated queries in which the tool invocation steps are often explicitly hinted at. Moreover, each query requires multiple steps to resolve, necessitating the model to plan the sequence of tool invocations.
@@
    Real deployed tools. We provide an evaluation platform deployed with tools across various categories, such as perception, operation, logic, and creativity. All tools are executable rather than simulated by text description. A detailed and executable ground truth tool chain is provided for each task, including each tool-use step and the final answer. Each step includes the tool name, argument value, and return value. The detailed tool chains enable a fine-grained evaluation of the actual problem-solving abilities of tool agents.
@@
Table 2: Comparison of GTA queries with AI-generated queries. The steps and tool types for queries in ToolBench and m&m’s are explicitly stated, as marked in red and blue. The queries in APIBench are simple and only contain one step. Our GTA’s queries are both step-implicit and tool-implicit.
@@
We manually design 229 real-world tasks and corresponding executable tool chains to evaluate mainstream LLMs. We build a platform covering a total of 14 tools across perception, operation, logic, and creation categories. Tools and some data samples are illustrated in Figure 1. We design fine-grained tool evaluation metrics that cover the entire process of tool invocation. Our findings indicate that real-world scenario queries present challenges to existing LLMs, with GPT-4 completing fewer than 50% of the tasks and most LLMs managing less than 25%.
@@
  2. An evaluation platform equipped with a wide variety of executable tools covering the categories of perception, operation, logic, and creativity. Fine-grained metrics are designed for tool use, unveiling tool-augmented LLMs’ reasoning and planning capabilities in real-world scenarios.
@@ 2 Related Work
Tool-use evaluations. With the rise of LLM-based agents, many studies have been conducted to evaluate the tool-use capabilities of LLMs. ToolBench [23] collects RESTful APIs and leverages ChatGPT [1] to design tool-use tasks and corresponding tool chains. Two metrics, Pass Rate and Win Rate, are devised to evaluate the efficacy of tool use. APIBench [21] is a comprehensive dataset that includes APIs from HuggingFace, TorchHub, and TensorHub, with evaluation metrics focusing on Abstract Syntax Tree (AST) accuracy. API-Bank [12] comprises 53 commonly utilized APIs, such as SearchEngine, PlayMusic, BookHotel, and ImageCaption, along with a comprehensive tool-augmented LLM workflow to evaluate the API calling, retrieving, and planning abilities. m&m’s [14] is a benchmark to evaluate tool use for multi-step multimodal tasks. It aims to evaluate different planning strategies for LLMs as planning agents. Most of the benchmarks above, however, rely on AI-generated queries. The tool-use steps are explicitly and rigidly included. Thus, these queries do not accurately represent real-world scenarios. Among many previous studies, GAIA [16] is renowned for its real-world scenario based benchmark aiming at evaluating general AI assistants, which is closer to our work. It designs conceptually simple questions for humans yet is challenging for most advanced AIs. However, GAIA focuses on artificial general intelligence (AGI). In contrast, GTA is designed to evaluate tool agents specifically, offering real-deployed tools and executable tool chains for a fine-grained evaluation in real-world scenarios. Osworld [30] is also a real-world benchmark featuring multi-step, complex tasks inspired by authentic user cases. Still, it is specifically tailored for computer environments, whereas GTA is devised for tool agents operating in more generalized real-world scenarios.
@@ 3.1 Dataset Formulation
Given a set of tools $\mathcal{T}_{c}=\left\{t_{k}\right\}_{k=1}^{N}$, a sample in GTA is composed of five parts $(\mathcal{F},\mathcal{Q},\mathcal{T},\mathcal{C},\mathcal{A})$. Among these parts, $\mathcal{F}$ is a set of files containing one or two images. $\mathcal{Q}$ is a query based on $\mathcal{F}$. It is a real-world scenario based problem of simple form but needs to be solved through multiple steps with tools in $\mathcal{T}_{c}$. Which tools need to be used, and in what steps are not explicitly included in the query. They require reasoning and planning by the LLM, which serves as a central controller. This procedure is given in the reference tool chain $\mathcal{C}=\{s_{i}\}_{i=1}^{m}$. The tool chain contains $m$ steps. Each step is $s_{i}=(t_{i},a_{i},r_{i})$, where $t_{i}$ is the tool used in step $i$. $a_{i}$ and $r_{i}$ indicate arguments and return values. $\mathcal{T}=\bigcup_{j=1}^{m}\{t_{j}\}\subseteq\mathcal{T}_{c}$ notes the set of tools involved in this query. $\mathcal{A}$ is the final answer yielded by the LLM after reasoning with tools.
In our setting, $\mathcal{T}_{c}$ contains 14 tools across four categories, including perception, operation, logic, and creativity. The full list of tools is shown in Figure 1, and more detailed information can be found in Appendix B.1. The queries $\mathcal{Q}$ are classified into three types: subjective, objective, and image generation. Examples of the three types of queries are shown in Appendix B.2. For a subjective query $\mathcal{Q}_{s}$, the final answer $\mathcal{A}$ is usually some descriptive text. It is not unique, but the general idea is the same. In this case, $\mathcal{A}$ contains a list of three reference answers. For an objective query $\mathcal{Q}_{o}$, $\mathcal{A}$ is a uniquely determined number or phrase. For an image generation query $\mathcal{Q}_{g}$, we do not measure the generated image directly. In this situation, $\mathcal{A}=\emptyset$.
@@ 3.2 Query Construction
To construct $(\mathcal{F},\mathcal{Q},\mathcal{T})$, we first gather human-designed queries that meet three main principles: i) Given $\mathcal{T}\subseteq\mathcal{T}_{c}$, the task $(\mathcal{F},\mathcal{Q})$ can be solved with the capabilities enabled by tools in $\mathcal{T}$. ii) To evaluate LLMs’ reasoning and planning abilities, the tool invocation steps should not be explicitly stated in the queries. iii) The queries are meaningful and based on real-world scenarios. Satisfying all the principles simultaneously is challenging. It requires $\mathcal{F},\mathcal{Q}$, and $\mathcal{T}$ to match each other sensibly and logically. We use a query construction pipeline based on exemplar expansion, as shown in the first part of Figure 2. We first give some initial exemplars with diverse scenarios and tool combinations. Then, we instruct annotators to create more queries based on the exemplars.
@@
Exemplar designed by experts. We first design some initial questions as exemplars, which are provided in Appendix C.1. These example questions are of diverse scenarios and contain different tool combinations. Every sample should comprise six components: $\mathcal{F}$ (image files), $\mathcal{Q}$ (queries), $\mathcal{T}$ (involved tools), $\mathcal{S}$ (solution steps), $\mathcal{A}$ (answers), and $\mathcal{E}$ (evidence). Image files $\mathcal{F}$ could be obtained from the internet, and their URLs must be recorded. $\mathcal{F}$ could also be a photo taken or a diagram drawn by the annotators. The query $\mathcal{Q}$ must avoid obvious references to a specific tool. For example, the query please describe the image for me is unqualified since it obviously refers to the tool ImageDescription. The components $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{E}$ will not appear in the final dataset but are utilized to assist annotators in meeting the annotation requirements. $\mathcal{S}$ represents the steps required to solve the problem. Annotators should note down the steps, ensuring their number exceeds two. The answer $\mathcal{A}$ of objective queries should be given to guarantee a unique answer. To ensure the uniqueness, the answer should not depend on the images generated in previous steps. For example, the question what kind of animal is in the picture should not be asked after generate an image of an animal, as the answer is uncertain. For queries utilizing the Google Search tool, $\mathcal{E}$ should include the answer’s URL and a screenshot pinpointing the answer’s location to verify the query’s searchability with the tool.
@@
Diversified expansion by annotators. After the initial exemplars are given, we instruct annotators to create more samples based on each exemplar. We adopt a diversified expansion strategy for the annotators to expand the questions based on the exemplars. The general idea is to keep the tool set $\mathcal{T}$ of the template unchanged or slightly modify it. Then, annotators brainstorm scenarios different from the template. Further information on the diversified expansion approach is detailed in Appendix C.2. For each sample, we have crafted a manual expansion example to serve as guidance for the annotators. After the expansion process, we perform a quality check and manually filter out the questions that do not satisfy the expansion requirements. The instruction documents for annotators are reported in Appendix C.3.
@@
Considerations for the search tool. Web search tools like Google Search may return variable results over time, harming the accuracy and stability of evaluation. To address this problem, we perform two constraints. First, the question must be answered using web search rather than relying on an LLM’s internal knowledge. This can be achieved by designing time-sensitive questions, such as what is the 2023 QS ranking of Tsinghua University rather than general inquiries like where is Tsinghua University located in China. We may also direct the query to a specific information source, for instance, asking what is the recipe for Ma Po Tofu according to the BBC Good Food website instead of a broad question like what is the recipe for Ma Po Tofu. Second, it is crucial to ensure that answers remain constant over time within an evaluation dataset. To fulfill this criterion, we can specify a time frame, web page, or organization within the question. An example would be what is the 2024 QS ranking of Tsinghua University, rather than what is the QS ranking of Tsinghua University.
@@ 3.3 Tool Chain Construction
Based on the $(\mathcal{F},\mathcal{Q},\mathcal{T})$ samples constructed in Section 3.2, we instruct three annotators majoring in computer science to manually construct the corresponding tool chain $\mathcal{C}$ and the final answer $\mathcal{A}$. We design a JSON file structure containing the query-related tool list, image paths, and ReAct [33] style dialog sequences. The dialog sequences include the user query, the executable tool chain, and the final answer. Initially, $(\mathcal{T},\mathcal{F},\mathcal{Q})$ are put into the associated sections for tools, images, and user queries. Subsequently, we deploy all tools in $\mathcal{T}_{c}$. The annotators utilize the tools according to the reference steps $\mathcal{S}$ and get the outcomes. They record this process in the tool chain section of the dialog sequences, alongside the final answer. Since we do not evaluate the tools’ efficacy, when a tool fails to provide accurate recognition for a query (for instance, OCR inaccuracies in text recognition within diagrams), we discard the query. Through the above process, we ensure the feasibility of the questions, the executability of the tool chains, as well as the precision of the final answers. The structure of the tool chain is provided in Appendix C.4.
@@
Total tool calls $557$